OpenStudy (anonymous):

cos(2 tan−1 x) simplify the expression

OpenStudy (campbell_st):

I think you need a triangle and use \[\tan^{1} (x)\] to find the missing side. tan = opp/adj so x/1 gives opp = x and adj = 1  so now you have \[\cos(2 \theta) = \cos^2(\theta) - \sin^2(\theta)\] looking at the triangle \[\sin(\theta) = \frac{x}{\sqrt{x^2 +1}} ...and.... \cos(\theta) = \frac{1}{\sqrt{x^2 + 1}}\] so your equation now becomes \[\cos(2 \theta) = (\frac{1}{\sqrt{x^2 + 1}})^2 - (\frac{x}{\sqrt{x^2 + 1}})^2\] you just need to simplify the equation there is a common denominator is \[x^2 + 1\]